OpenStudy (anonymous):

A money lender borrows some money at 4% per annum and pays the interest at the end of the year. He lends it at 6% p.a. compounded half-yearly and receives the interest at the end of the year. In this way he gains Rs.104.50 during the year. How much money does he borrow?

ganeshie8 (ganeshie8):

lets say he borrowed \(x\) dollars

OpenStudy (anonymous):

x*4*1/100-x(1+6/2/100)=104.50

ganeshie8 (ganeshie8):

very good try ! keep in mind : `compounded half-yearly` means, n = 2, right ?

OpenStudy (anonymous):

yes so only rate of interest =6/2 and time =2 yrs but didnt put that one

ganeshie8 (ganeshie8):

interest paid = 4% of x = \(\large \dfrac{4}{100}x\) interest earned = 6% compounded half yearly = \(\large x\left(1+\dfrac{0.06}{2}\right)^2 - x\)

ganeshie8 (ganeshie8):

subtracting above both interests gives you the profit earned

ganeshie8 (ganeshie8):

Profit = \[\large x\left(1+\dfrac{0.06}{2}\right)^2 - x - \large \dfrac{4}{100}x\]

ganeshie8 (ganeshie8):

set it equal to 104.50 and solve x

OpenStudy (anonymous):

3.2036x=104.50 so x=32.069

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

x=32.65

OpenStudy (anonymous):

4.2406x-x-0.04

ganeshie8 (ganeshie8):

check ur calculation for first term : \(\large x\left(1+\dfrac{0.06}{2}\right)^2\)

OpenStudy (anonymous):

oh i made a mistake by not dividing it by 2 ya

OpenStudy (anonymous):

2.06*2.06 frankly i wrote

ganeshie8 (ganeshie8):

okay, so what do you get for x ?

OpenStudy (anonymous):

94.97 something is it right?

ganeshie8 (ganeshie8):

nope, try again

OpenStudy (anonymous):

0.0209x

OpenStudy (anonymous):

5000

ganeshie8 (ganeshie8):

\[\large x\left(1+\dfrac{0.06}{2}\right)^2 - x - \large \dfrac{4}{100}x\] \[\large 1.0609x - x - 0.04x\] \[\large 0.0209x\]

OpenStudy (anonymous):

so 5000

ganeshie8 (ganeshie8):

yes, setting it equal to 104.50 and solving x gives you x = 5000 !